4. Suppose that z = f(x,y) and x = y2 + and y = 1 - 2st. Use the chain rule to find af/as and affat
Answer #1

Given z=f(x,y) and a=52t2 and y=1-25t use chain rule. ole e (64+2 ds And y = (1-25+) P = -24 put in formula of = .(25) + 4.(-Put in formula. JA - .(2+) + 3y 1-281 | Answers